OpenStudy (tylerascough):

epress 59/100 as a decimal

OpenStudy (tylerascough):

express sorry

OpenStudy (anonymous):

anything over 100, has 0. in the front: 27/100 =0.27 89/100 =0.89 48/100 =0.48

OpenStudy (anonymous):

so the answer would be 0.59

OpenStudy (owen_turner):

Divide 59 by 100 then you will get your answer.